OpenStudy (anonymous):

62) If angle 1 and angle 2 form linear pair such the measure of angle 1 = 3x+5deg. and the measure of angle 2 = 4x+35deg., find the measures of both angles. a) 45 degrees & 135 degrees b) 55 degrees & 125 degrees c) 65 degrees & 115 degrees d) 75 degrees & 105 degrees

OpenStudy (cwrw238):

i've never heard of a 'linear pair' is it this: 

Parth (parthkohli):

A linear pair is two adjacent angles that sum up to 180, so the equation is: \(\Large \color{purple}{\rightarrow 3x + 5 + 4x + 35 = 180 }\) \(\Large \color{purple}{\rightarrow 7x + 40 = 180 }\) \(\Large \color{purple}{\rightarrow 7x = 140 }\) \(\Large \color{purple}{\rightarrow x = 20 }\) So, angle 1 is 3(20) + 5 = 65 and angle 2 is 4(20) + 35 = 115

Parth (parthkohli):

@cwrw238 Sure, two supplementary adjacent angles form a linear pair.

Parth (parthkohli):

You get the answer as the option c.

OpenStudy (cwrw238):

right- thank you - never heard of that in uk schools